A 1-kg ball moving at a velocity of 12m/s collides head-on with a 2-kg ball moving at 24 m/s in the opposite direction. Calculate the velocity of each ball after collision if (a) e=2/3, (b) the balls stick together and (c) the collision is perfectly elastic.


1
Expert's answer
2021-03-18T14:05:12-0400

Mass of 1st ball(m1) = 1 kg

Velocity of 1st ball("u_1") = 12 m/s

Mass of 2nd ball(m2) = 2 kg

Velocity of 2nd ball("u_2") = -24m/s

Coefficient of restitution e = "\dfrac{2}{3}"

(a) Velocity of 1st ball(v1) = "\dfrac{(m_1-em_2)u_1+(1+e)m_2u_2}{m_1+m_2}=-28m/s"

Velocity of 2nd ball(v2) = "\dfrac{(m_2-em_1)u_2+(1+e)m_1u_1}{m_1+m_2}=-4m/s"

(b) If the balls stick with each other, they both move with the same final velocity(V) after collision

"V=\dfrac{m_1u_1+m_2u_2}{m_1+m_2}=-12m/s"

(c) For perfectly elastic collision, e = 1,

"v_1=\dfrac{(m_1-m_2)u_1+2m_2u_2}{m_1+m_2}=-36m/s"

"v_2=\dfrac{(m_2-m_1)u_2+2m_1u_1}{m_1+m_2}=0m/s"